OpenStudy (anonymous):

8E+05 * 3E+13 can someone please help me understand how to solve this? :)

OpenStudy (jdoe0001):

bear in mind that Engineering notation is written like so E+05 really means \(\bf 10^5\) E+13 really means \(\bf10^{13}\) so 8E+05 really means \(\bf 8 \times 10^5\)

OpenStudy (jdoe0001):

and also bear in mind that \(\bf 10^5 = 10 \times 10 \times 10 \times 10 \times 10 = 100000\) \(\bf 10^5\) is 10 and 5 zeros

OpenStudy (anonymous):

use order of operations

OpenStudy (anonymous):

Thanks jdoe0001

OpenStudy (jdoe0001):

yw